when you divide Brian's farther age by 2,3,4,5 or 6 the remainder is 1. if Brian's faryher is less than 100 years how old is he
Found 2 solutions by Theo, ikleyn:
Answer by Theo(13342)

least common multiple of 2, 3, 4, 5, 6 is 60, i believe.
60/2 = 30
60/3 = 20
60/4 = 15
60/5 = 12
60/6 = 10
add one to 60 and you should get your answer.
61/2 = 30 with a remainder of 1.
61/3 = 20 with a remainder of 1.
61/4 = 15 with a remainder of 1.
61/5 = 12 with a remainder of 1.
61/6 = 10 with a remainder of 1.


Answer by ikleyn(52806)